40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 309 of 336  Not logged in ELOG logo
ID Date Author Type Category Subject
  1410   Thu Mar 19 10:45:43 2009 YoichiConfigurationIOOA loose wire found for MC1
I attached a 6-hour trend of the MC mirror OSEM signals with the MC unlocked.
The drift of the MC1 is within 20 counts (0.6um in terms of each OSEM).
This is comparable to the other MC mirrors.
Attachment 1: AfterWireFix-1.pdf
AfterWireFix-1.pdf
  1409   Thu Mar 19 02:45:36 2009 YoichiConfigurationIOOA loose wire found for MC1
I found a loose connection of a wire in the cross-connect between an ADC and the MC1 coil driver's UL bias input.
I tightened it.
To see if this fixes the MC1 drift problem, I will do another round of MC1 drift measurement.
You can lock the MC if you need to use the IFO but please note it in the elog.

Thanks.
  1408   Tue Mar 17 08:44:37 2009 YoichiConfigurationIOOMC1 drift
I'm done with the MC1 drift measurement.
The result is attached. It is clear that MC1 is in trouble. The small drifts in the MC2/MC3 are insignificant compared to the crazy MC1 behavior.
Since there is no drift in the coil feedback voltage monitors, it is probably not a problem of the DACs.
We may be able to fix this by pushing the cables for the MC1 satellite amplifier. But it may require replacement of the coil driver.


Quote:
There seems to be a large drift of MC1 even when there is no WFS feedback.
The attached plot is an example a 20min trend. You can see that MC1 OSEM signals drift significantly larger than that of MC2/MC3.
You can also be sure that there is no drifting voltage applied to the coils on the MC1 during this period.

If no one is working on the IFO today during the LV meeting, I'd like to leave the MC unlocked and see the trend of the MC1 OSEM signals.
Please do not turn on the MC auto locker unless you want to use the IFO.
If you want to do some measurements, please go ahead and lock the MC, but please write it down in the elog.
Thanks.
Attachment 1: MC1_Drift3.pdf
MC1_Drift3.pdf
  1407   Mon Mar 16 15:19:52 2009 OsamuDAQElectronicsSR785

I borrowed SR785 to measure AA, AI noise and TF.

  1406   Mon Mar 16 12:26:59 2009 YoichiConfigurationIOOMC1 drift
There seems to be a large drift of MC1 even when there is no WFS feedback.
The attached plot is an example a 20min trend. You can see that MC1 OSEM signals drift significantly larger than that of MC2/MC3.
You can also be sure that there is no drifting voltage applied to the coils on the MC1 during this period.

If no one is working on the IFO today during the LV meeting, I'd like to leave the MC unlocked and see the trend of the MC1 OSEM signals.
Please do not turn on the MC auto locker unless you want to use the IFO.
If you want to do some measurements, please go ahead and lock the MC, but please write it down in the elog.
Thanks.
Attachment 1: MC1_Drift1.pdf
MC1_Drift1.pdf
  1405   Mon Mar 16 01:20:40 2009 ranaConfigurationIOOMCWFS noise filtered on the SUS-MC
Recently, we noticed that the IOO-WFS system runs at 2048 Hz and sends its signals to the MC SUS
systems which run at 16 kHz. There is no upsampling filter or anti-imaging filter.

So, I've implemented an RLP666 filter as FM1 in the SUS-MCn_ASC(PIT/YAW) filter banks. This is like a 4th order
Cheby low pass with a low Q notch at 2048 Hz to catch the first image.

The attached PNG shows the ASCPIT_OUT signals before and after the filter is implemented. As you can see, the
big aliased spikes are gone. The reason that MC2 is different from MC1/3 is that they have a hardware 28Hz low pass
and MC2 doesn't. So MC2 had a 28 Hz low pass in software already to match the actuation phase between all the MC
mirrors. The apparent power law noise floor from 40-300 Hz in MC2 is not real - just the Hanning window tail.

And yes, it has been this way for several years and none of us noticed. It remains to be seen if this was causing
any noise in the MC coil drivers via slew rate limiting.
Attachment 1: xarm.png
xarm.png
  1404   Sun Mar 15 21:50:29 2009 Kakeru, Kiwamu, OsamuUpdateComputersSome computers are rebooted

We found c1lsc, c1iscex, c1iscey, c1susvme, c1asc and c1sosvme are dead.
We  turned off all watchdogs and turned off all lock of suspensions.
Then, I tried to reboot these machines from terminal, but I couldn't login to all of these machines.

So, we turned off and on key switches of these machines physically, and login to them to run startup scripts.

Then we turned on all watchdogs and restored all IFO.

Now they look like they are working fine.
 

  1403   Sat Mar 14 22:53:12 2009 KakeruUpdateoplevsarm cavity oplev calibration
I finished a calibration of optical levers.

To calibrate oplevs, I locked appropriate cavity and tilted a mirror.
A cavity with tilted mirror decrease its arm power. So I can know how much the tilt is.
For calibration of ITMX and ETMX, I locked X arm and measured TRX.
For ETMX, ETMY and BS, I locked Y arm and measured TRY
For PRM, I locked PRC and measured SPOB
For SRM, I locked SRC and measured REFL166

I used, for example, C1:SUS-ITMX_OPLEV_PERROR as an oplev signal.

The calibration factors for each mirror is below. The attachment is figures of my fitting.
I used modified equation for ITM calibration from my last calibration, so the value become small around 30%.

ITMX Pitch: 142   microrad/counts
ITMX Yaw:   145   microrad/counts
ITMY Pitch: 257   microrad/counts
ITMY Yaw:   206   microrad/counts

ETMX Pitch: 318   microrad/counts
ETMX Yaw:   291   microrad/counts
ETMY Pitch: 309   microrad/counts
ETMY Yaw:   299   microrad/counts

BS Pitch:    70.9 microrad/counts
BS Yaw:      96.3 microrad/counts

PRM Pitch:   78.5 microrad/counts
PRM Yaw:     79.9 microrad/counts

SRM Pitch:  191   microrad/counts
SRM Yaw:    146   microrad/counts


It looks strange that ITMY, BS and SRM has different value. I think this is a fitting problem.
These data have some asymmetry and cause these 20%-30% difference.
Actually, PRM Yaw has a little asymmetry but the value doesn't differ from Pitch.
This means that this calibration factor potentially has below 30% error.
(These data are the most fine data. I think we must adjust Y arm yaw alignment. The beam spot of ETMY looks too low!)
For SRM, I couldn't get fine data because it was very sensitive to tilt and easily lose its lock.
When I tuned cavity enough, The data become almost flat, so I used detuned cavity.

It is also strange that ITMX and ITMY is different. I guess that this is caused by the difference of the QPD input. The sum of QPD is around 10000 for ITMX and around 4500 for ITMY.
The difference between BS or PRM and SRM is same, I guess. The sum of QPD input for BS and SRM is around 1500, but for SRM, it is around 10000.

I will write more detailed document and upload it with my calibration code.
Attachment 1: oplev.pdf
oplev.pdf oplev.pdf oplev.pdf
  1402   Fri Mar 13 22:07:14 2009 YoichiUpdateLockingCalibrated XARM error signal spectrum
Of course I made a mistake.
I put a pole at 1525Hz whereas it should have been a zero.

The correct calibration factor is:
G: 4.2e-13
P:
Z: 1525

I attached a revised spectrum.


Quote:
I did a rough calibration of the XARM error spectrum.
See the attached calibrated spectrum.

I started from this Rana's elog entry.
http://www.ldas-sw.ligo.caltech.edu/ilog/pub/ilog.cgi?group=40m&task=view&date_to_view=04/07/2005&anchor_to_scroll_to=2005:04:07:20:28:36-rana

I first injected a 20Hz sin signal into C1:SUS-ETMX_LSC_EXC and measured the response to the ETMX SUSPOS.
Using the calibration of the SUSPOS given in the above entry, I calibrated the ETMX coil actuation efficiency.
It was 3.4e-12 m/cnt @20Hz for C1:SUS-ETMX_LSC_EXC.

Then I locked the X-arm and injected a calibration peak at 20Hz.
From the ratio of the peaks in C1:SUS-ETMX_LSC_IN2 and C1:LSC-XARM_IN1, I calibrated the X-arm error signal to be 4.2e-13 m/cnt.
We have to also take into account the cavity pole of the arm, 1525Hz (the design value, may not be actual).
So I used the following calibration in the DTT:

G: 4.2e-13
P: 1525
Z:

Note that the attached spectrum shows the actual motion of the X-arm (or equivalent frequency noise) after suppressed by the feedback servo,
unlike conventional noise spectra showing "virtual" displacement which would have been induced in the absence of servos.
Attachment 1: XarmErrorSpeCalibrated.pdf
XarmErrorSpeCalibrated.pdf
  1401   Fri Mar 13 20:23:37 2009 YoichiUpdateLSCAO path transfer function with X-arm locked
I measured the AO path transfer function while the X-arm is locked with the POX PDH signal.
The POX-I signal was already connected to the input 1 of the CM board. So I injected a signal from the EXC-B channel of the board and measured the transfer function from TP2B to TP1A. To open the loop, I disabled the switch befor the EXC-B.
The attached plot shows the measured transfer function.
There is a bump around 2kHz, which can also be seen in the AO path TF posted in elog:1399, but not the large structure at around 3.8kHz.
The 3.8kHz structure is probably created by the feedback.
Attachment 1: AOPath-Xarm.png
AOPath-Xarm.png
  1400   Fri Mar 13 19:26:09 2009 YoichiUpdateDMFseisBLRMS compiled

 I compiled seisBLRMS.

The tricks were the following:

(1) Don't add path in a deployed command.
It does not make sense to add paths in a compiled command because it may be moved to anywhere. Moreover, it can cause some weird side effects. Therefore, I enclosed the addpath part of mdv_config.m in a "if ~isdeployed ... end" clause to avoid adding paths when deployed. Instead of adding paths in the code, we have to add paths to necessary files with -I options at the compilation time. This way, mcc will add all the necessary files into the CTF archive.

(2) Add mex files to the CTF archive by -a options.
For some reason, mcc does not add necessary mex files into the CTF archive even though those files are called in the m-file which is being compiled. We have to add those files by -a options.

(3) NDS_GetData() is slow for nodus when compiled.
NDS_GetData(), which is called by get_data() stops for a few minutes when using nodus as an NDS server.
This problem does not happen when not compiled. I don't know the reason. To avoid this, I modified seisBLRMS.m so that when an environmental variable $NDS is defined, it will use an NDS server defined in this variable.

I wrote a Makefile to compile seisBLRMS. You can read the file to see the details of the tricks.
I also wrote a script start_seisBLRMS, which can be found in /cvs/cds/caltech/apps/DMF/compiled_matlab/seisblrms/. To start seisBLRMS, you can just call this script.
At this moment, seisBLRMS is running on megatron. Let's see if it continues to run without crashing.

Quote:

The seisBLRMS has been running on megatron via an open terminal ssh'd into there from allegra with matlab running. This

is because I couldn't get the compiled matlab functionality to work.

Even so, this running script has been dying lately because of some bogus 'NDS' error. So for today I

have set the NDS server for mDV on megatron to be fb40m:8088 instead of nodus.ligo.caltech.edu. If this seems to fix the problem

I will make this permanent by putting in a case statement to check whether or not the mDV'ing machine is a 40m-martian or not.

  1399   Fri Mar 13 05:16:21 2009 YoichiUpdateLockingLocking update
Yoichi, Osamu,

With adjustments of the loop gains during the CARM offset reduction, the IFO reaches arm_power = 25 sort of robustly unless the 3.8kHz oscillation rings up.
At arm_power = 25, the CARM and DARM start to oscillate at around 400Hz. Probably I need more gain tweaks.
Annoying thing is that the 3.8kHz oscillation sometimes rings up suddenly and kills the lock.
This can happen anywhere above arm_power = 6 or so.
Because of a strange structure in the CARM loop gain around 3.8kHz, we cannot increase the CARM UGF beyond 1kHz.
The attached plots are the AO path open loop transfer function (attm2 is the zoom of attm1) measured at arm_power = 13.

Tomorrow, I will lock the X-arm and measure the transfer function from the AO path input to the X-arm error signal to see
if there is the same structure at 3.8kHz (X-arm error signal has the 3.8kHz peak).
Attachment 1: AOTF2.png
AOTF2.png
Attachment 2: AOTF2-zoom.png
AOTF2-zoom.png
  1398   Thu Mar 12 20:59:04 2009 KakeruUpdateIOOMC drift is terrible
After Rana went for his dinner, I aligned periscope to make the MC output 3.2 (Attachment 1).

After that, to align WFS, I unlocked the MC, unlocked the MZ and decrease the beam power to WFS QPD, and re-centerd WFC beam.
I restored MZ and MC lock.
I enabled MC autolocker, and change C1:IOO-WFS_Gain_Slider from 0 to 0.02 to lock WFS.


Quote:
Kakeru, Rana, Yoichi

We used the SUS DRIFT MON screen to set the MC biases such that the mirrors were returned to the old OSEM values.
To do this, we set the nominals and tolerances using the appropriate scripts in the mDV/extra/C1/ directory.

We then used the MC_ALIGN screen to set the angle bias sliders.

Then Kakeru and I went to the PSL table to the periscope magic and maximize the MC transmission. Kakeru seems to
have the careful Japanese alignment touch and I am hungry, so I am leaving him to optimize the power. After he
finishes he is going to align the beam to the WFS and turn the MC autolocker back on. The x-arm is locked on a
TEM00 mode so the MC alignment is maybe OK.
Attachment 1: MCtrans090312.png
MCtrans090312.png
  1397   Thu Mar 12 19:11:27 2009 ranaUpdateIOOMC drift is terrible
Kakeru, Rana, Yoichi

We used the SUS DRIFT MON screen to set the MC biases such that the mirrors were returned to the old OSEM values.
To do this, we set the nominals and tolerances using the appropriate scripts in the mDV/extra/C1/ directory.

We then used the MC_ALIGN screen to set the angle bias sliders.

Then Kakeru and I went to the PSL table to the periscope magic and maximize the MC transmission. Kakeru seems to
have the careful Japanese alignment touch and I am hungry, so I am leaving him to optimize the power. After he
finishes he is going to align the beam to the WFS and turn the MC autolocker back on. The x-arm is locked on a
TEM00 mode so the MC alignment is maybe OK.
  1396   Thu Mar 12 18:48:37 2009 YoichiUpdateIOOMC aligned but ...
After the MZ alignment, I aligned the MC with the periscope mirrors.
It looked like the MC mis-alignment was mainly caused by the input beam change.
So I left the MC mirrors as they were to keep the output beam pointing.
However, after I finished the alignment, the MC output beam was too low on the Faraday.
Also the X-arm did not lock to TEM00 mode. So the MC mirrors must have also shifted to a weird alignment state.
I should have restored the MC mirror alignment to a good state using the OSEM DC signals.

Rana came in and restored the MC mirror alignment using the SUS drift mon.
He and Kakeru is now working on the periscope to align the beam into the MC.
  1395   Thu Mar 12 18:44:02 2009 YoichiUpdatePSLMZ aligned
The MC lost the alignment somehow this afternoon.
So I thought it was good time to touch the MZ because I had to align the MC using the periscope anyway.

I mainly touched the mirror with a PZT. The MZ reflection went down from 0.5 to 0.3.
  1394   Thu Mar 12 15:57:53 2009 YoichiUpdateIOOMC drift is terrible
Yoichi, Osamu,

Last night's locking work was totally interrupted by the sabotage by the MC.

First, after I measured the RF_AM, the MC alignment was somehow shifted largely and the MC did not lock to TEM00 mode.
I only mis-aligned MC2 to measure the RF_AM, but the MC reflection beam was also shifted (looking at the WFS QPD), that means MC1 was mis-aligned somehow.
Moreover, even when the MC is not locked, i.e. no feedback to the mirrors, the OSEM values of the MC mirrors (all of them) drift a lot in 10min scale.
I was totally puzzled. So I rebooted c1iovme and c1sosvme. Then this strange drift of the OSEM values stopped.
Even though, the MC tended to lose lock within ten minutes because the WFS QPDs were not centered.
We did several iterations of re-centering and finally the MC started to stay locked happily. The MC reflection beam was symmetric.

Then this morning when I came in (to be honest, afternoon), the MC reflection looked asymmetric again. The WFS QPDs were mis-centered again.
The attached files show an 8-hour trend of various MC related signals.
There was a half-degree temperature change starting from around 11AM. Corresponding to that, the IOO-QPD signals drifted indicating that the PSL beam pointing
was shifted. The MZ PZT signal shows a similar trend, so the beam pointing may have been shifted by the MZ (not sure).
The MC WFS, transmission QPD signals show the same trend.
This is too bad.

Right now, the PSL beam pointing is monitored by the QPDs detecting the transmitted beam through the first mirror of the periscope.
This means even if we can track the beam pointing drift with the QPDs, we can't correct the beam pointing using the periscope mirrors.
I don't want to touch the MZ mirrors for this purpose.
I propose to put a pick-off mirror after the second mirror of the periscope to send light to the IOO-QPDs. This way, we can use the periscope
mirrors to restore the beam pointing screwed up by the MZ.
Attachment 1: MC_Drift-1.pdf
MC_Drift-1.pdf
Attachment 2: MC_Drift-2.pdf
MC_Drift-2.pdf
  1393   Thu Mar 12 02:18:42 2009 YoichiUpdateLockingMC_I spectra (RF_AM)
I took several spectra of MC_I signal (see attm1).

The blue curve is when the MC was locked. The green curve (RF_AM) shows the MC_I spectrum when the MC is unlocked and MC2 is mis-aligned,
so that no resonance should happen. The brown curve is when the PSL shutter was closed (dark noise).
There are some structures in the green curve but not at 3.8kHz.

The second attachment compares the MC_I spectrum (the same as the green one in the first attachment) with the Xarm error signal.
Of course these two spectra were taken at different times.

Some of the peaks in the X-arm error signal seem to be coming from the MC RF_AM.
Attachment 1: MC_I_Spe.png
MC_I_Spe.png
Attachment 2: MC_I-Xarm.png
MC_I-Xarm.png
  1392   Thu Mar 12 00:29:39 2009 JenneOmnistructureDMFDMF being whiny again

Quote:
The seisBLRMS has been running on megatron via an open terminal ssh'd into there from allegra with matlab running.


[Yoichi, Jenne]

seisBLRMS was down again. I assumed it was just because the DMF Master Enable was in the 'Disabled' state, but enabling it didn't do the trick. Rana's green terminal window was complaining about not being able to find nodus.ligo.caltech.edu. Yoichi and I stopped it, closed and restarted Matlab, ran mdv_config, then ran seisBLRMS again, and it seems happy now.

On the todo list still is making the DMF / seisBLRMS stuff happy all the time.
  1391   Wed Mar 11 23:41:33 2009 Kakeru, YoichiUpdateIOOWFS centering

We found the MC reflection was distorted . And WFC beam went to upward of QPD

We recentered WFC beam and these problems were fixed

  1390   Wed Mar 11 22:57:48 2009 YoichiUpdateLockingCalibrated XARM error signal spectrum
I did a rough calibration of the XARM error spectrum.
See the attached calibrated spectrum.

I started from this Rana's elog entry.
http://www.ldas-sw.ligo.caltech.edu/ilog/pub/ilog.cgi?group=40m&task=view&date_to_view=04/07/2005&anchor_to_scroll_to=2005:04:07:20:28:36-rana

I first injected a 20Hz sin signal into C1:SUS-ETMX_LSC_EXC and measured the response to the ETMX SUSPOS.
Using the calibration of the SUSPOS given in the above entry, I calibrated the ETMX coil actuation efficiency.
It was 3.4e-12 m/cnt @20Hz for C1:SUS-ETMX_LSC_EXC.

Then I locked the X-arm and injected a calibration peak at 20Hz.
From the ratio of the peaks in C1:SUS-ETMX_LSC_IN2 and C1:LSC-XARM_IN1, I calibrated the X-arm error signal to be 4.2e-13 m/cnt.
We have to also take into account the cavity pole of the arm, 1525Hz (the design value, may not be actual).
So I used the following calibration in the DTT:

G: 4.2e-13
P: 1525
Z:

Note that the attached spectrum shows the actual motion of the X-arm (or equivalent frequency noise) after suppressed by the feedback servo,
unlike conventional noise spectra showing "virtual" displacement which would have been induced in the absence of servos.
Attachment 1: XarmErrorSpeCalibrated.pdf
XarmErrorSpeCalibrated.pdf
  1389   Wed Mar 11 21:03:51 2009 Kakeru and KiwamuUpdateIOOPSL angle QPD

Kakeru and Kiwamu

We placed a QPD on the PSL bench for PSL angle monitor.

 

Quote:

I checked a broken QPD, which was placed for PSL angle monitor, and finally I cocluded one segment of the quadrant diode was broken.

The broken segment has a offset voltage of -0.7V after 1st I-V amplifier. It means the diode segment has a current offset without any injection of light.

Tomorrow I will check a new QPD for replacement.

Kiwamu IZUMI

 

 As we mentioned before, old QPD which used to be placed is broken.

 And we put broken QPD into the "photodiodes" box under the soldering table.

 

 

  1388   Wed Mar 11 16:53:48 2009 YoichiUpdateLockingJunks in around kHz
Rana, Yoichi

Last night, we tried to find out the source of the kHz region peaks in the DARM and CARM error signals.
These peaks are also present in the error signal of the single arm locking by RF (both X and Y).
The attachment 1 shows spectra of MC_F and XARM error signal when XARM is locked by the POX PDH signal.
There is a sharp peak at 3.8kHz in MC_F. This peak was there in a reference spectrum taken on June 24 2008.

In the XARM error signal, there is also a broad peak around 3.8kHz. This peak moves between 3.75kHz and 3.8kHz from time to time.
(the brown curve was taken when the peak moved to 3.75kHz).
Also there is a notch like structure at 3.8kHz in the XARM error spectrum. Looks like the peak in the MC_F is creating a notch here, but
no idea why.

We tapped on the PSL table, the end chambers and the SPOB table and looked at the spectra to see if there is any change.
Rana also developed a cool Walkie-Talkie excitation technique, where he put one of the walkie-talkies on the PSL table by the MZ and yelled at the other one while looking at a DTT screen in the control room.
None of these had any effect on the XARM error, while MC_F responded to the disturbances.

We also turned on and off the steering mirror PZT closed loop buttons, moved the PMC, MZ and the ISS gain sliders and changed the MC gain, offset.
Nothing affected the XARM error.

Osamu found old spectra of the XARM signal (attm2). The legends say DARM but these are XARM signals.
Almost the same structures can be seen including the notch at 3.8kHz. Seems like it's been like this for long time.

We should check, RF-AM, MC coil dirivers, Piezo-Jena noise etc.
Attachment 1: MC_F-XARM.pdf
MC_F-XARM.pdf
Attachment 2: old-xarm.pdf
old-xarm.pdf
  1387   Wed Mar 11 16:41:22 2009 steveUpdateMOPAspare NPRO power

The spare M126N-1064-700,  sn 5519 of Dec 2006 rebuilt  NPRO's power output

 measured   750mW at DC2.06A with Ohpir meter.

Alberto's controller  unit 125/126-OPN-PS,  sn516m was disconnected from lenght measurment NPRO on the AP table.

5519 NPRO  was clamp to the optical table  without heatsink and it was on for 15 minutes.

  1386   Wed Mar 11 14:51:01 2009 Kakeru, Joe, RobUpdateIOOMC alignment

This morning, MC alignment was gone and MC wasn't lock.

We checked old value of pitch, yaw, and position offset of each MC mirror, and found they were jumped.

We don't know the reason of this jump, but we restore each offset value and MC backed to lock.

  1385   Wed Mar 11 11:30:15 2009 josephbConfigurationCameras 

I modified the Video.db file used by c1aux located in  /cvs/cds/caltech/target/c1aux.

I added the following channels to the db file, intended for either read in or read out by the digital camera scripts.

C1:VID-ETMY_X_COM
C1:VID-ETMY_Y_COM

C1:VID-ETMY_X_STDEV

C1:VID-ETMY_Y_STDEV

C1:VID-ETMY_XY_COVAR

C1:VID-ETMY_EXPOSURE

C1:VID-ETMY_GAIN

C1:VID-ETMY_X_UL

C1:VID-ETMY_Y_UL

C1:VID_ETMY_X_SIZE

C1:VID_ETMY_Y_SIZE

A better naming scheme can probably be devised, but these will do for now.

  1384   Wed Mar 11 04:33:57 2009 ranaConfigurationComputer Scripts / Programswild ndsproxy tclshexe

The ndsproxy tcl task on nodus was eating up all the CPU and making the elog slow. I killed it and restarted it.

It looks like it hasn't been making a log file since January. Someone who has some skill in decoding the cryptic csh stdout redirection

syntax should look into this (its in target/ndsproxy/)

  1383   Wed Mar 11 01:16:40 2009 ranaSummaryIOOrogue trianglewave in the MC Servo offset slider

On Monday evening, I ran this command: trianglewave C1:IOO-MC_REFL_OFFSET 0 4 120 600;ezcawrite C1:IOO-MC_REFL_OFFSET 1.76

which I thought (from the syntax help) would move that offset slider with a period of 120 seconds for 600 seconds. In actuality, the last argument is the

run time in number of periods. So the offset slider has been changing by 8 Vpp for most of the last day. Oops. The attached image shows what effect

this had in the MC transmitted power (not negligible). This would also make the locking pretty difficult.

 

In the second plot you can see the zoom in view for ~30 minutes. During the first part, the MCWFS are on and there are large fluctuations

in the transmitted power as the WFS offset changes. This implies that the large TEM00 carrier offset we induce with the slider couples into

the WFS signals because of imbalances in the quadrant gains - we need someone to balance the RF gains in the WFS quadrants by injecting

an AM laser signal and adjusting the digital gains.

 

Since there is still a modulation of the MC RFPD DC with the WFS on, we can use this to optimize the REFL OFFSET slider. The third plot

shows a 8 minute second trend of this. Looks like the slider offset of zero would be pretty good.

Attachment 1: Untitled.png
Untitled.png
Attachment 2: Untitled.png
Untitled.png
Attachment 3: a.png
a.png
  1382   Tue Mar 10 04:55:41 2009 YoichiUpdateLockingLocking: 3.7kHz large oscillation
Yoichi, Jenne, Alberto,

As I reported on the last Thursday, there is a large oscillation in CARM and DARM error signals (attm1).
I put notch filters (3.75kHz, Q=10, 30dB) in the CARM and DARM loops. This let us go up to the arm power of more than 20 and stay there for a while.
The dashed curves in the attm1 are the spectra when the notches are off, and the solid curves are when the notches are used.
We could somewhat suppress the DARM peaks but not CARM.
Of course this is clearly not a good solution. We should find the cause of the oscillation and kill it.

Attm2 is the spectrum of the PO_DC signal flowing in the CM board measured by the SR785. More specifically, CH1 is TP1A and CH2 is TP2A of the CM board.
This was taken right after the AO path was engaged. At this stage, the AO path gain is very low. But you can already see a seed of the oscillation in the spectrum.

Attm3 shows the same spectra taken after the arm power is increased to 4 but before the PO_DC hand off. You can see large peaks around 3.75kHz.
After this, the peaks grow as the power goes up.

Attm4 is the loop gain of the AO path after the PO_DC hand off (arm power = 4).
Attm5 is the zoom of the same TF around 3.7kHz. Clearly there is something wrong at this frequency. We should check the CM board and the MC board as well as the SPOB PD.

One time I was able to go up to arm power = 27 or so. At this power level, the DARM loop started to oscillate, probably, around the UGF.
However because of the 3.7kHz problem, we can't stay at this power level long enough to make diagnostic measurements (like open loop TF).
We should tackle the 3.7kHz issue first.
Attachment 1: CARM_DARM_Spectra.pdf
CARM_DARM_Spectra.pdf
Attachment 2: PODC_Spe_AOPath_Engaged.png
PODC_Spe_AOPath_Engaged.png
Attachment 3: PODC_Spe_before_PODC_handoff.png
PODC_Spe_before_PODC_handoff.png
Attachment 4: AOGain3.png
AOGain3.png
Attachment 5: AOGain2.png
AOGain2.png
  1381   Mon Mar 9 23:55:38 2009 OsamuDAQComputersbscteststand and kami1 outside martian

This morning there was a confliction of tpman running on fb40m and kami1. Alex fixed it temporary but Rana suggested it was better to move both PCs outside martian. We moved both PCs physically to the control room and connected to general network with a local router. I believe it won't conflict anymore but if you guess these PC might have trouble please feel free to shutdown.

 

Today's work summary:

 *connected expansion chassis to bscteststand

 *obtained signals on dataviewer, dtt for both realtime and past data on bscteststand with 64kHz timing signal

 

Questions:

Excitation channels are not shown, only "other" is shown.

qts.mdl should run with 16kHz but 16kHz timing causes a slow speed on dataviewer and failing data aquisition on dtt. We are using 64kHz timing but is it really correct?

  1380   Mon Mar 9 23:13:22 2009 YoichiUpdateComputer Scripts / Programstdsdata doesn't work

Quote:

We confirmed that new tds(/cvs/cds/caltech/apps/linux/tds_090304/) works well on linux 64, and replaced it to /cvs/cds/caltech/apps/linux/tds/

The old /cvs/cds/caltech/apps/linux/tds is put in /cvs/cds/caltech/apps/linux/tds.bak

 The tdscntr.pl in the new tds was probably the one from LLO, which is actually the version I sent to Tobin. It had paths and channel names defined for the LLO. So I copied back my original 40m version.

  1379   Mon Mar 9 19:33:10 2009 ranaUpdateDMFseisBLRMS in temp condition

The seisBLRMS has been running on megatron via an open terminal ssh'd into there from allegra with matlab running. This

is because I couldn't get the compiled matlab functionality to work.

Even so, this running script has been dying lately because of some bogus 'NDS' error. So for today I

have set the NDS server for mDV on megatron to be fb40m:8088 instead of nodus.ligo.caltech.edu. If this seems to fix the problem

I will make this permanent by putting in a case statement to check whether or not the mDV'ing machine is a 40m-martian or not.

Attachment 1: Untitled.png
Untitled.png
  1378   Mon Mar 9 19:27:16 2009 ranaConfigurationComputersMove of the CLIO Digital Controls test setup

Because of the network interference we've had from the CLIO system for the past 3-4 days, I asked the guys to remove

the test stand from the 40m lab area. It is now in the 40m control room. Since it needed an ethernet connection to get out

for some reason we've let them hook into GC. Also, instead of using a real timing signal slaved to the GPS, Jay suggested

just skipping it and having the Timing Slave talk to itself by looping back the fiber with the timing signal. Osamu will enter

more details, but this is just to give a status update.

  1377   Mon Mar 9 17:11:38 2009 AlbertoConfigurationoplevsoptical levers centering

Yoichi, Alberto

this afternoon we centered the optical levers for all the optics.

To do that we first ran the alignment scripts for all the cavities.

  1376   Mon Mar 9 16:54:52 2009 Kakeru, RanaUpdateComputer Scripts / Programstdsdata doesn't work

We confirmed that new tds(/cvs/cds/caltech/apps/linux/tds_090304/) works well on linux 64, and replaced it to /cvs/cds/caltech/apps/linux/tds/

The old /cvs/cds/caltech/apps/linux/tds is put in /cvs/cds/caltech/apps/linux/tds.bak

  1375   Mon Mar 9 14:57:30 2009 KakeruUpdateComputer Scripts / Programstdsdata doesn't work

I tested new tdsdata and found it was working well.

I excited C1:SUS-ITMY_SUSPIT_EXC with tdssine, and get data from C1:LSC-TRY_OUT (testpoint) and C1:SUS- ITMY_OPLEV_PERROR (recorded point) with new and old tdsdata.

With old tdsdata (/cvs/cds/caltech/apps/linux/tds/bin/tdsdata), I found some jumps of datapoint, which is a same problem with before (Attachment 1).

With new tdsdata (/cvs/cds/caltech/apps/linux/tds_090304/bin/tdsdata), there looks to be no jumps (Attachment 2; taken about 10 minutes after Attachment 1).

The problem of old tdsdata looks to be remaining even for recordedpoints.

You should use /cvs/cds/caltech/apps/linux/tds_090304/bin/tdsdata.

Quote:

Matt logged in and rebuilt the TDS stuff for us on Mafalda in /cvs/cds/caltech/apps/linux/tds_090304.

He says that he can't build his stuff on 64-bit because there's not a sanctioned 64-bit build of GDS yet.

This should have all the latest fixes in it. I tried using both the old and new code from allegra and they both are fine:

./tdsdata 16384 2 C1:IOO-MC_F > /users/rana/test.txt

I loaded the data I got with the above command and there were no data dropouts. Possibly the dropout problem is only

associated with testpoints and so we have to wait for the TP fix.

 

Attachment 1: oldtds.png
oldtds.png
Attachment 2: newtds.png
newtds.png
  1374   Mon Mar 9 12:04:18 2009 YoichiUpdateComputersTPs and AWG are back
I had to do one more reboot of tpman and daqd to get the TPs working.
I confirmed the alignment scripts run fine.

Now the oplevs of some optics are largely mis-centered. Alberto and I will center them after lunch.
  1373   Mon Mar 9 11:09:33 2009 AlbertoUpdateComputersRe: Not even data retrieval working

Quote:
Although getting the regular DAQ data works, we can't get any testpoints.

I tried restarting tpman several times; there's no inittab on fb40m for this so we should get Alex to set one up when he comes.
I also tried various power cycles and reboots: daqawg, daqctrl, etc. I also notice that Osamu's setup of new stuff is connected to
the same rack and power strips as all of our sensitive DAQ machines. We should find out if there was any hardware installed in the
last couple days; it would be easy to accidentally unplug or damage on of our fibers.

I moved the old tpman.log over to tpman.log.090308. It starts out with a header and then just lists when each TP is requested.

When restarting tpman it puts the following into the terminal:
fb:controls>./tpman &
[1] 1037
fb:controls>VMIC RFM 5565 (0) found, mapped at 0x2868c90
VMIC RFM 5579 (1) found, mapped at 0x2868c90
Could not open 5565 reflective memory in /dev/daqd-rfm1
16 kHz system
Spawn testpoint manager
Channel list length for node 0 is 4168
Test point manager (31001001 / 1): node 0
which is OK?; its the same startup outputs that are in the old log file. It would be nice if there was not and error message about the RFM.
Requesting new testpoints via tdsdata, dtt, or the diag command line doesn't seem to work. tpman doesn't spit anything out although 'tp show 0'
does show that the TP is selected.

Once Alex fixes the 'tpman' issue, we should make sure to put an inittab or startup script in there so that tpman writes a log
file and also archives its old log files upon a restart.


Alex fixed the problem. It was caused by the awgtpman running on kami1.martian which conflicted with the tpman in fb0.

Killing awgtpman on kami1 allowed for the tpman on tp0 to work properly again.

If more test points are needed, Alex suggested to tune the GDS settings accordingly.
What this actually means, I still have to understand it.
  1372   Mon Mar 9 10:59:05 2009 AlanOmnistructureComputersssh agent on fb40m restarted for backup

After the boot-fest, the nightly backup to Powell-Booth failed, and an automatic email got sent to me. I restarted the ssh agent, following the instructions in /cvs/cds/caltech/scripts/backup/000README.txt .

  1371   Sun Mar 8 23:14:52 2009 ranaUpdateComputer Scripts / Programstdsdata doesn't work

Matt logged in and rebuilt the TDS stuff for us on Mafalda in /cvs/cds/caltech/apps/linux/tds_090304.

He says that he can't build his stuff on 64-bit because there's not a sanctioned 64-bit build of GDS yet.

This should have all the latest fixes in it. I tried using both the old and new code from allegra and they both are fine:

./tdsdata 16384 2 C1:IOO-MC_F > /users/rana/test.txt

I loaded the data I got with the above command and there were no data dropouts. Possibly the dropout problem is only

associated with testpoints and so we have to wait for the TP fix.

  1370   Sun Mar 8 23:09:26 2009 ranaUpdateComputersNot even data retrieval working
Although getting the regular DAQ data works, we can't get any testpoints.

I tried restarting tpman several times; there's no inittab on fb40m for this so we should get Alex to set one up when he comes.
I also tried various power cycles and reboots: daqawg, daqctrl, etc. I also notice that Osamu's setup of new stuff is connected to
the same rack and power strips as all of our sensitive DAQ machines. We should find out if there was any hardware installed in the
last couple days; it would be easy to accidentally unplug or damage on of our fibers.

I moved the old tpman.log over to tpman.log.090308. It starts out with a header and then just lists when each TP is requested.

When restarting tpman it puts the following into the terminal:
fb:controls>./tpman &
[1] 1037
fb:controls>VMIC RFM 5565 (0) found, mapped at 0x2868c90
VMIC RFM 5579 (1) found, mapped at 0x2868c90
Could not open 5565 reflective memory in /dev/daqd-rfm1
16 kHz system
Spawn testpoint manager
Channel list length for node 0 is 4168
Test point manager (31001001 / 1): node 0
which is OK?; its the same startup outputs that are in the old log file. It would be nice if there was not and error message about the RFM.
Requesting new testpoints via tdsdata, dtt, or the diag command line doesn't seem to work. tpman doesn't spit anything out although 'tp show 0'
does show that the TP is selected.

Once Alex fixes the 'tpman' issue, we should make sure to put an inittab or startup script in there so that tpman writes a log
file and also archives its old log files upon a restart.
  1369   Sat Mar 7 16:50:25 2009 YoichiUpdateComputersNot even data retrieval working
Now our digital system is really in trouble.
We can't even get data from tp channels.

I did another round of computer reboots, this time including the RFM bypass switch, c0daqctrl, c0dcu1 and fb40m itself.
But the problem still persists.

I guess there is nothing I can do until Alex comes in.
  1368   Fri Mar 6 18:26:37 2009 YoichiConfigurationComputersezca tools and tds tools work around
I updated the wrapper scripts so that they do not retry more than 6 times.
Otherwise, the wrapper scripts loop over infinitely when you give wrong arguments.



Quote:
Some of ezca commands and tds commands sporadically fail with a segmentation fault on linux machines.
As far as I know, ezcawrite, ezcastep, ezcaswitch, and tdswrite have this problem.
These are commands to write values into epics channels. So usually people do not check the exit status of those commands in their scripts.
This could cause incomplete execution of, for example, down scripts.
Ideally, this problem should be fixed in the source codes of the problematic commands.
However, I don't have a patience to wait it to happen, and I needed to fix these problems immediately for the lock acquisition.
So I resorted to a hacky solution.
I renamed those commands to *.bin, e.g. ezcawrite -> ezcawrite.bin.
Then wrote wrapper scripts to repeatedly call those commands until it succeeds.
For example, ezcawrite now looks like,
#!/bin/csh -f
setenv POSIXLY_CORRECT
while (! { ezcawrite.bin $* })
      echo "Retry $0 $*"
end
So, when ezcawrite.bin fails, the command retries it and show a message "Retry ....".

If you need to call the original commands, you can always do so by adding ".bin" at the end of the command name.
Currently the following commands are wrapped.
ezcawrite, ezcaservo, ezcastep, ezcaswitch, tdswrite, tdssine.

Please let me know if you have any trouble with this.
  1367   Fri Mar 6 18:22:42 2009 YoichiSummaryComputersScripts to restart the FE computers
While doing locking, the FE computers are overloaded sometimes and I have to reboot them.
Being sick of logging into the FE computers one by one to start front end codes, I wrote scripts to do this automatically.
The scripts are in /cvs/cds/caltech/scripts/FE/.
For example, you can restart c1lsc by typing
restartFE c1lsc
You can give multiple computer names to the restartFE command like,
restartFE c1lsc c1asc c1susvme1

To restart all the FE computers, type
restartFE all

For the scripts to work properly, the computers have to accept login, i.e. you either have to power cycle the computers or push "Reset" buttons on the RFMNETWORK medm screen prior to running the scripts.
  1366   Fri Mar 6 18:14:58 2009 YoichiUpdateComputersawg not working
Starting from this afternoon, the awg is not working.
I rebooted FE computers, c0daqawg as well as tpman and daqd processes on fb40m several times.
But the problem is still there.
I sent an email to Alex.
  1365   Fri Mar 6 15:23:39 2009 YoichiUpdateLockingLocking distracted by the QPD whitenning problem again
By looking at the time series of DARM signal at the time of a lock loss, the oscillation frequency was about 3.5kHz (see the attm1 and its zoomed version attm2).
I will measure the DARM loop gain around this frequency next.


Quote:
Tonight, I was able to ramp up the arm power to around 20. Then the DARM loop started to oscillate and the IFO lost lock in a few seconds.
I repeated this several times, then realized that the transmission QPDs were not working properly again due to the well known sticky slider problem.
I should have run slider_twiddle script. Since the DARM RF signal is normalized by the sqrt(TRX+TRY), it is reasonable that the DARM loop got unstable.

The fact that I was able to go up to arm power = 20 means there is nothing saturating below this power level.
Attachment 1: lockLoss3.pdf
lockLoss3.pdf
Attachment 2: lockLoss3-zoom.pdf
lockLoss3-zoom.pdf
  1364   Fri Mar 6 05:44:14 2009 YoichiUpdateLockingLocking distracted by the QPD whitenning problem again
Tonight, I was able to ramp up the arm power to around 20. Then the DARM loop started to oscillate and the IFO lost lock in a few seconds.
I repeated this several times, then realized that the transmission QPDs were not working properly again due to the well known sticky slider problem.
I should have run slider_twiddle script. Since the DARM RF signal is normalized by the sqrt(TRX+TRY), it is reasonable that the DARM loop got unstable.

The fact that I was able to go up to arm power = 20 means there is nothing saturating below this power level.
  1363   Fri Mar 6 01:04:49 2009 Kiwamu IZUMIConfigurationIOO!! lock-in amp disconnected !!

The power supply of a lock-in amp, which is on the Y-arm side of PSL clean room, was pulled out by my mistake.

Then I reconnected it, but I don't know whether it is re-adjusted properly.

I'm sorry about this. If you are using that amp, it should be checked.

  1362   Thu Mar 5 23:18:38 2009 KakeruConfigurationComputerstdsdata doesn't work
I found that tdsdata doesn't work.

When I star tdsdata, he takes a few ~ 10 seconds of data, and he dies with a message "Segmentation fault".
I tried to get data for some times and some channels, and this problem was observed everytime.
I also tried tdsdata on allegra, op440m and mafalda, and it didn't work on all of them.

Yesterday, I got a new version of tdsdata (which modified the problem of Message ID: 1328) and tried to build 
thme on my directory (/cvs/cds/caltech/users/kakeru.....)
This may have some relation to this problem.
  1361   Thu Mar 5 05:07:09 2009 YoichiUpdateLockingWednesday night locking
Tonight, I was having a problem with the PO_DC hand-off.
It fails most of the time.
I increased the averaging time for the PD1_DC offset measurement.
I also wrote a script to match the gain of the transmission DC and the PO_DC signals.
This script (/cvs/cds/caltech/scripts/CM/matchPODCGain ) measures the gains of the old (TRX+TRY) and new (PO_DC) signals at 150Hz and returns the optimal value to be put into the input matrix.
cm_step script calls matchPODCGain to determine the matrix element value for the PO_DC signal.

Even with this script, the hand-off was still unreliable.
I checked the AO path loop gain just before the hand off. It looked normal.
Then I realized that the oscilloscope I hooked up to the PO_DC signal using a T-BNC may be introducing some noise into the channel.
So I removed it. Then the PO_DC hand off went well at least once.
The IFO still loses lock at around arm power 10.

I attached time series of the latest lock loss. The second attachment is a zoom of the first one.
This time, there is a glitch in the ETM feedback signals, which is also present in the DARM and CARM and error signals.
I saw this kind of glitches several times today.
Attachment 1: lockLoss5.png
lockLoss5.png
Attachment 2: lockLoss5-zoom.png
lockLoss5-zoom.png
ELOG V3.1.3-